# UniTable: Towards a Unified Framework for Table Structure Recognition via Self-Supervised Pretraining

### v1

@@ 7 Conclusion
We present UniTable, a training framework that unifies both the training paradigm and training objective of TSR. Its training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by SSP from diverse unannotated tabular images. Our framework unifies the training objectives of all three TSR tasks, extracting table structure, cell content, and cell bbox, into a unified task-agnostic training objective: language modeling. Extensive quantitative and qualitative analyses highlights UniTable’s SOTA performance on four of the largest TSR datasets To promote reproducible research, enhance transparency, and SOTA innovations, we open-source our code and release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TSR datasets, supporting all three TSR tasks.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning, especially TSR. To enable users to easily try UniTable on their own tabular images and obtain fully digitized HTML tables, we release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TSR datasets, supporting all three TSR tasks.
Appendix A Using UniTable in Practice
@@
Appendix B Effectiveness and Scalability of SSP
Table 4 demonstrates the effectiveness and scalability of SSP on all four subsets of SynthTabNet. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Pretraining the visual encoder on 1M tabular images provides a significant improvement, and pretraining on 2M images continues to increase the performance on all TSR tasks.
Appendix C Generalization of the Unified Training Objective
UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TSR. Table 5 shows results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS and table cell bbox detection evaluated with mAP.
Appendix D Visualization of the Visual Codebook
Fig. 4 and 5 present the original table overlaid by the selected token indices from the 2M VQ-VAE codebook. We highlight the token indices under investigation in red and use blue for the others. The input image size is $448\times 448$ and the patch size is $16\times 16$, so each image has $28\times 28$ tokens chosen from the 16384 entries in the codebook. First, we highlight the tokens representing the blank white background. Comparing the red and blue indices, we observe that the red region has covered most of the blank space inside the table, and the blue region has formulated a convex hull tightly surrounding either the texts or separation lines. Next, we highlight the tokens representing the separation lines. Comparing the red and blue regions, we observe that the red regions align with all the separations within the table, e.g., line separations and color shading separations. Taking a closer look at the indices, we find the codebook has learned to represent an abstract concept by assigning multiple tokens, where each reflects a different scenario. For example, token “111” represents the separation above the gray color shading and token “964” represents the one below, and token “15282” represents the separation above a bold horizontal line and “10807” represents the one below. A certain type of separation also has multiple choices depending on the portion of the line within the patch, e.g., token “14181”, “8714”, and “10807” all representing the “below a bold horizontal line”, but the differences lie in the line thickness and amount of black inside the image patch. Such a fine-grained categorization also explains why the codebook needs so many tokens to represent the implicit conventions created by humans in the table.
@@ Appendix D Visualization of the Visual Codebook
Table 5: UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TSR. Results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS and cell bbox detection with mAP.

### v2

@@ 6 Conclusion
We present UniTable, a training framework that unifies both the training paradigm and training objective of TR. Its training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by SSP from diverse unannotated tabular images. Our framework unifies the training objectives of all three TR tasks, extracting table structure, cell content, and cell bbox, into a unified task-agnostic training objective: language modeling. Extensive quantitative and qualitative analyses highlights UniTable’s SOTA performance on four of the largest TR datasets To promote reproducible research, enhance transparency, and SOTA innovations, we open-source our code and release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TR datasets, supporting all three TR tasks.
Appendix A Detail comparisons with milestone VLMs
Table 4: UniTable outperforms GPT-4V and LLaVA-v1.6 by a huge margin for table recognition.
Model Finance PubTabNet Marketing Sparse
Samples 100 300 500 100 300 500 100 300 500 100 300 500
LLaVA-v1.6-vicuna-7b 37.72 35.57 34.16 41.66 42.19 41.80 36.54 34.34 34.14 36.95 37.42 38.49
LLaVA-v1.6-mistral-7b 38.40 39.76 38.07 39.16 37.79 38.96 34.16 33.20 32.25 40.28 35.97 37.06
LLaVA-v1.6-vicuna-13b 43.50 45.13 43.95 48.17 49.05 49.68 41.37 42.00 41.40 46.38 46.04 47.21
LLaVA-v1.6-34b 40.86 42.96 40.69 46.30 47.99 49.30 35.08 35.75 33.99 39.13 39.60 39.85
GPT-4-1106-vision-pre. 64.05 66.93 66.13 69.26 69.86 69.62 58.51 60.03 59.13 55.64 57.02 56.99
UniTable Large 99.63 99.56 99.57 99.54 99.54 99.55 99.11 99.10 99.10 99.34 99.35 99.34
Table 4 shows detailed comparisons (S-TEDS score) with GPT-4V and the latest LLaVA v1.6. In summary, UniTable outperforms GPT-4V and LLaVA-v1.6 by a huge margin for table recognition.
Dataset: SynthTabNet [28], is a large-scale table benchmark that offers control over dataset size, table structure, style, and content type. It has 4 subsets: Finance, PubTabNet, Marketing, Sparse. We incrementally evaluate more samples up to 500 on each subset because no further notable changes in the results are observed.
Models: We compare UniTable with GPT-4V and LLaVA. For GPT-4V, we use the gpt-4-1106-vision-preview. For LLaVA, we test on all 4 variants of its latest version, v1.6, provided by the official model weights on HuggingFace [21]: vicuna-7b, mistral-7b, vicuna-13b, 34b. We use the same prompt as in the latest work that evaluates GPT-4V’s OCR capabilities by Shi et al., [35]: “Please read the table in this image and return an HTML-style reconstructed table in text, do not omit anything.” The max generated sequence length is set to 1500 to cover the largest table in the test.
Appendix B Using UniTable in Practice
@@
Appendix C Effectiveness and Scalability of SSP
Table 10 demonstrates the effectiveness and scalability of SSP on all four subsets of SynthTabNet. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Pretraining the visual encoder on 1M tabular images provides a significant improvement, and pretraining on 2M images continues to increase the performance on all TR tasks.
Appendix D Generalization of the Unified Training Objective
UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TR. Table 11 shows results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS and table cell bbox detection evaluated with mAP.
Appendix E Visualization of the Visual Codebook
Fig. 5 and 6 present the original table overlaid by the selected token indices from the 2M VQ-VAE codebook. We highlight the token indices under investigation in red and use blue for the others. The input image size is $448\times 448$ and the patch size is $16\times 16$, so each image has $28\times 28$ tokens chosen from the 16384 entries in the codebook. First, we highlight the tokens representing the blank white background. Comparing the red and blue indices, we observe that the red region has covered most of the blank space inside the table, and the blue region has formulated a convex hull tightly surrounding either the texts or separation lines. Next, we highlight the tokens representing the separation lines. Comparing the red and blue regions, we observe that the red regions align with all the separations within the table, e.g., line separations and color shading separations. Taking a closer look at the indices, we find the codebook has learned to represent an abstract concept by assigning multiple tokens, where each reflects a different scenario. For example, token “111” represents the separation above the gray color shading and token “964” represents the one below, and token “15282” represents the separation above a bold horizontal line and “10807” represents the one below. A certain type of separation also has multiple choices depending on the portion of the line within the patch, e.g., token “14181”, “8714”, and “10807” all representing the “below a bold horizontal line”, but the differences lie in the line thickness and amount of black inside the image patch. Such a fine-grained categorization also explains why the codebook needs so many tokens to represent the implicit conventions created by humans in the table.
@@ Appendix E Visualization of the Visual Codebook
Appendix F Discussions on TSR Dataset Annotations
Word-wise vs. cell-wise bbox annotations. Fig. 7(a) is the word-wise annotation from table “PMC1574335 table 0” in PubTables-1M and Fig. 7(b) is the cell-wise annotation from table “PMC5897438 004 00” in PubTabNet. In cell-wise annotation, each cell has a unique bbox that can be matched to the non-empty cells in table structure HTML tags. However, in word-wise annotation, a bbox is matched to a single word, which cannot be easily combined with the table structure as in cell-wise annotation, which limits its general applicability.
((a)) Word-wise bbox from PubTables-1M
((b)) Cell-wise bbox from PubTabNet
Figure 7: Word-wise vs. cell-wise bbox annotations: (a) is the word-wise annotation from table “PMC1574335 table 0” in PubTables-1M and (b) is the cell-wise annotation from table “PMC5897438 004 00” in PubTabNet. In cell-wise annotation, each cell has a unique bbox that can be matched to the non-empty cells in table structure HTML tags. However, in word-wise annotation, a bbox is matched to a single word, which cannot be easily combined with the table structure as in cell-wise annotation, which limits its general applicability.
F.1 Discussions on TSR Dataset Annotations
((a)) Unrelated text around table
((b)) Overlapping bbox
Figure 8: Our UniTable has successfully identified inconsistent table cell bbox annotations in PubTables-1M, which can guide future data annotation pipelines and qualities.
Our goal is to push forward the development of the document and table understanding, thus we hope to see high-quality and consistent data annotations. Because of the unified language modeling framework formulation, we discover previously unacknowledged inconsistency in one of the largest TR datasets, PubTables-1M.
Here, we have identified three types of inconsistent table cell bbox annotations: (1) Fig. 8(a) (“PMC4964067 table 7”) shows bbox annotations for unrelated text around table; (2) Fig. 8(b) (“PMC6202420 table 1”) shows overlapping bbox annotations; (3) Bbox annotations exceed the boundary of image size, e.g., [-4.6, 278.6, 19.5, 292.4] from “PMC4802837 table 1”. With a rapid filter comparing the image size with the bbox, we find that a surprising 53.10% (402914 over 758849) of the table annotations in the training set have at least one bbox going beyond the image size. We hope that these inconsistencies discovered by UniTable can help guide future data annotation pipelines and qualities.
Appendix G SOTA results
Table 5: Comparisons on IC19B2M
Model IoU 0.6 WAvg. F1
NLPR-PAL [8] 30.50 20.60
CascadeTabNet [31] 35.40 23.20
GTE [41] 38.50 24.80
UniTable Base (Ours) 54.97 40.15
UniTable Large (Ours) 58.10 42.62
Table 6: Comparisons on PubTabNet
Model S-TEDS TEDS $\text{AP}_{50}$ $\text{AP}_{75}$
EDD [42] 89.90 88.30 79.20 -
GTE [41] 93.01 - - -
TableFormer [28] 96.75 93.60 82.10 -
TableMaster [40] 96.04 96.16 - -
OTSL [25] 95.50 - - 88.00
VAST [12] 97.23 96.31 94.80 -
UniTable Base (Ours) 95.63 94.78 97.94 87.27
UniTable Large (Ours) 97.89 96.50 98.43 92.44
Table 7: Comparisons on SynthTabNet
Model S-TEDS $\text{AP}_{50}$ $\text{AP}_{75}$
TableFormer [28] 96.70 87.70 -
DRCC [34] 98.70 - -
UniTable Base (Ours) 98.97 98.99 98.79
UniTable Large (Ours) 99.39 99.00 98.87
Table 8: Comparisons on FinTabNet
Model S-TEDS
GTE [41] 91.02
EDD [42] 90.60
OTSL [25] 95.90
TableFormer [28] 96.80
VAST [12] 98.63
UniTable Base (Ours) 97.19
UniTable Large (Ours) 98.89
Table 9: Comparisons on PubTables-1M
Model $\text{AP}_{50}$ $\text{AP}_{75}$
Faster R-CNN [36] 81.50 78.50
OTSL [25] - 89.60
DETR [36] 97.10 94.80
UniTable Base (Ours) 94.48 88.64
UniTable Large (Ours) 95.68 93.28
Table 5, 6, 7, 8, and 9 show detailed comparisons across five of the largest table benchmarks from our paper. In summary, UniTable outperforms prior methods and achieves SOTA on four out of the five largest table datasets. (“-” means not reported by the compared methods.)
@@ Appendix G SOTA results
Table 11: UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TR. Results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS and cell bbox detection with mAP.